# On Dirichlet Series Involving $\zeta(s)$ and Extensions of the Euler–Mascheroni Constant

### v1

@@ 3. Further Proposition and Examples
For $n,m\in\mathbb{N}$, and $\mu\in\mathbb{N}_{0}$
$\begin{array}[]{l}\lim\limits_{\mu\rightarrow\infty}{\gamma}(n,m;\mu)=\begin{cases}\zeta(m),&\text{if}\;\;m>1,\\[1.0pt] \lim\limits_{\mu\rightarrow\infty}H_{\mu},&\text{if}\;\;m=1,\\ \end{cases}\\[19.0pt] \lim\limits_{n\rightarrow\infty}{\gamma}(n,m;\mu)=\begin{cases}H_{\mu}^{(m)},&\quad\;\text{if}\;\;\mu>1,\\ 0,&\quad\;\text{if}\;\;\mu=0,\\ \end{cases}\\[18.0pt] \lim\limits_{m\rightarrow\infty}{\gamma}(n,m;\mu)=\begin{cases}1,&\qquad\;\;\,\text{if}\;\;\mu>1,\\ 0,&\qquad\;\;\,\text{if}\;\;\mu=0.\end{cases}\end{array}$
@@
First, we establish the convergence of $D(s_{1},s_{2};\lambda)$. For $k\geq 2$, we have
$0<\zeta(k)-1<\sum\limits_{m=1}^{\infty}\frac{2^{m}}{2^{mk}}\leq\frac{1}{2^{k-2}},$
@@
converges absolutely for all $s_{1},s_{2}\in\mathbb{C}$ and all $\lambda\in\mathbb{C}\setminus\mathbb{Z}_{<0}$ with $s_{1}\notin\mathbb{Z}_{\leq 0}$. Next, we derive an integral representation of $D(s_{1},s_{2};\lambda)$. To begin with, we prove the following lemma, which will be used in the subsequent analysis.
@@
Let $m\geq 2$. Then we have the following identity:
@@
More generally, for integers $m>k\geq 1$,
@@ 7. Functional relations
which converges absolutely for $|z|<1$, or for $|z|=1$ with $\Re(s)>1$, where we assume $\lambda\notin\mathbb{Z}_{\leq 0}$. By Definition $(1.1)$, we observe immediately that
@@
which holds for $\Re(s)<0$, $|z|<1$, and $0<\lambda\leq 1$ (see [Er, 1.11 (6)]), we obtain the desired functional relation. ∎

### v2

@@ 3. Further Proposition and Examples
For $n,m\in\mathbb{N}$ and $\mu\in\mathbb{N}_{0}$,
$\lim_{n\to\infty}\gamma(n,m;\mu)$ $=H_{\mu}^{(m)},$
[…]
@@
First, we establish the convergence of $D(s_{1},s_{2};\lambda)$. For $k\geqslant 2$, we have
$0<\zeta(k)-1<\sum\limits_{m=1}^{\infty}\frac{2^{m}}{2^{mk}}\leqslant\frac{1}{2^{k-2}},$
@@
converges absolutely for all $s_{1},s_{2}\in\mathbb{C}$ and all $\lambda\in\mathbb{C}\setminus\mathbb{Z}_{<0}$ with $s_{1}\notin\mathbb{Z}_{\leqslant 0}$. Next, we derive an integral representation of $D(s_{1},s_{2};\lambda)$. To begin with, we prove the following lemma, which will be used in the subsequent analysis.
@@
Let $m\geqslant 2$. Then we have the following identity:
@@
More generally, for integers $m>k\geqslant 1$,
@@ 7. Functional relations
which converges absolutely for $|z|<1$, or for $|z|=1$ with $\Re(s)>1$, where we assume $\lambda\notin\mathbb{Z}_{\leqslant 0}$. By Definition $(1.1)$, we observe immediately that
@@
which holds for $\Re(s)<0$, $|z|<1$, and $0<\lambda\leqslant 1$ (see [Er, 1.11 (6)]), we obtain the desired functional relation. ∎